News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 3 November 1999 9.00am EST

During the last 48 hours, Chandra completed observations of the GTO targets E0102-72.3, Pictor-A, NGC 2516, Hydra-A, RX-J0911.4+0551, IRAS 09104 and a series of calibration observations of NGC targets. In addition the recently discovered type 2 supernova in the galaxy NGC 1637 was successfully inserted into the schedule and observed as a Target of Opportunity.

The software patch to modify the on-board code to use the B-side LETG potentiometer value in place of the A-side value was uplinked and a test performed at 7am 10/3 to verify correct operation. The test was conducted during the pre-belt ACIS CTI measurements and verified that the correct value is being used by the software.

During the next 48 hours, Chandra will pass through the radiation belts, perform post-belt ACIS CTI measurements abd continue with ACIS observations of the Coma cluster, Mkn 421, Vega and GS 2000+25.

The spacecraft systems continued to operate nominally.

Return to Status Reports